# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 3. Quantum polymorphisms and reductions
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
@@
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
@@
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$.
We can then define a notion of q-definability, as follows.
@@
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{B}$ is q-definable from $\mathbb{A}$ if there exists a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
Observe, in ($(\mathbf{q_{1}})$), that a quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ is necessarily non-contextual, since all the vertices of $\mathbb{R}$ lie in a common constraint. (This will not be true anymore in the non-oracular setting, see Section 7.) Classical gadgets for CSPs that can be defined one from the other via pp-definitions provide logspace reductions between CSPs of different languages. The next result shows that the same is true for $\operatorname{CSP^{*}}$, provided that pp-definitions are replaced by the stronger q-definitions.
@@
Let $\mathbb{A}$ and $\mathbb{B}$ be structures such that $\mathbb{B}$ is q-definable from $\mathbb{A}$. Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{*}}(\mathbb{A})$ in logspace.
We prove Proposition 10 later in this section. For now, we focus on the following problem: How to construct q-definitions guaranteeing that a given $\operatorname{CSP^{*}}$ can be reduced to a different one? It is not hard to show that every q-definition yields a classical pp-definition. Indeed, considering ($(\mathbf{q_{1}})$) for quantum homomorphisms with a Hilbert space of dimension 1 (which corresponds to classical homomorphisms), we find that the tuples in $R$ are in one-to-one correspondence with the restrictions of homomorphisms $\mathbb{G}\to\mathbb{A}$ to $\{g_{1},\dots,g_{r}\}$. However, not every pp-definition is a q-definition. It turns out that, in order for a pp-definition to ensure the existence of a q-definition, we need to pair it with an extra construction making sure that the process of replacing constraints with gadgets preserves the existence of quantum homomorphisms. This is the notion of commutativity gadgets we define below, generalising [Ji].
Definition 11.
Fix a $\sigma$-structure $\mathbb{A}$. A commutativity gadget for $\mathbb{A}$ is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
  • $(\mathbf{c_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ extending $Q$.
  • $(\mathbf{c_{2}})$
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
Note that a commutativity gadget as in Definition 11 is not the same thing as a q-definition for $\mathbb{A}$ and $R=A^{2}$ in the sense of Definition 8.
Remark 12.
In [Zeman], an algebraic commutativity gadget is defined by replacing ($(\mathbf{c_{1}})$) by the following condition:
  • $(\mathbf{c_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
Note that this is a special case of $(\mathbf{c_{1}})$ and, thus, any commutativity gadget is an algebraic commutativity gadget. While this appears to be a strictly weaker requirement than the one in Definition 11, we shall prove below that the existence of an algebraic commutativity gadget in the sense of [Zeman] is equivalent to the existence of a commutativity gadget in our sense.
If $\mathbb{A}$ has a commutativity gadget, then we have a natural way of obtaining a sound reduction from $\operatorname{CSP^{*}}(\mathbb{B})$ to $\operatorname{CSP^{*}}(\mathbb{A})$ out of a classical gadget: For every pair of variables $x,y$ in a relation in an instance of $\operatorname{CSP^{*}}(\mathbb{B})$, glue to $x,y$ a copy of the commutativity gadget, identifying $x$ with $u$ and $y$ with $v$. The second condition in Definition 11 forces the associated projectors to commute in every quantum homomorphism to $\mathbb{A}$, and the first condition tells us that no additional constraint is forced upon $x,y$ by gluing the commutativity gadget. Formally, we prove the following result.
@@
Let $\mathbb{A},\mathbb{B}$ be structures such that $(i)$ $\mathbb{A}$ has a commutativity gadget, and $(ii)$ $\mathbb{B}$ is pp-definable from $\mathbb{A}$. Then $\mathbb{B}$ is q-definable from $\mathbb{A}$.
As a consequence of what we have discussed so far, the crucial step in order to quantise classical reductions from CSP theory is to build commutativity gadgets. The main result of this section classifies the languages for which commutativity gadgets exist, in terms of a property of their quantum polymorphisms, which we now define.
Definition 14.
A quantum polymorphism of arity $n\in\mathbb{N}$ for a structure $\mathbb{A}$ is a quantum homomorphism $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$.
We denote by $\operatorname{qPol}(\mathbb{A})$ the set of quantum polymorphisms of $\mathbb{A}$. Moreover, we define the quantum closure of $\operatorname{Pol}(\mathbb{A})$, denoted by $\operatorname{\oplus Pol}(\mathbb{A})$, as the union of the quantum closures of the set of $n$-ary polymorphisms $\operatorname{Hom}(\mathbb{A}^{n},\mathbb{A})$, for all $n\in\mathbb{N}$.
Remark 15.
Observe that $\operatorname{qPol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are both abstract minions (see the definition in [BBKO21, ciardo_quantum_minion]). Furthermore, it is not hard to check that $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are homomorphically equivalent, and $\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{A})$. In particular, it follows that, if $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$, then we have a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{A})$.
The main result of this section is the following.
Theorem 16.
Let $\mathbb{A}$ be a finite structure. The following are equivalent:
  1. (1)
    $\mathbb{A}$ has a commutativity gadget.
  2. (2)
    $\mathbb{A}$ has an algebraic commutativity gadget.
  3. (3)
  4. (4)
In the rest of this section, we provide the proofs of the results we stated.
Proof of Proposition 10.
For every symbol $R\in\rho$ of some arity $r$, let $\mathbb{G}_{R}$ be a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ (with its distinguished vertices $g_{1},\dots,g_{r}$). Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{*}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$. For $R\in\rho$ of some arity $r$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$, and we glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to_{q}\mathbb{A}^{\prime}$. Then, ($(\mathbf{q_{1}})$) guarantees the existence of a quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$ such that
$Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
for each $i\in[r]$ and each $a\in A$. Overall, this is a quantum homomorphism $\mathbb{Y}\to_{q}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\to_{q}\mathbb{A}$ is a quantum homomorphism. Consider the quantum function $Q^{\prime}\colon X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Observe that the restriction of $Q$ to $\mathbb{G}_{R}^{(\mathbf{x})}$ is a quantum homomorphism $\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$. By ($(\mathbf{q_{2}})$), this means that the assignment $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$). In particular, recalling that $x_{i}$ and $g_{i}^{(\mathbf{x})}$ are identified in $\mathbb{Y}$, it follows that the PVMs $Q^{\prime}_{x_{i}}=Q_{g_{i}^{(\mathbf{x})}}$ and $Q^{\prime}_{x_{j}}=Q_{g_{j}^{(\mathbf{x})}}$ commute for each choice of $i,j\in[r]$. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
(2) $\prod_{i\in[r]}Q^{\prime}_{x_{i},b_{i}}=0,$
which means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. ∎
Proof of Proposition 13.
Let $S=R^{\mathbb{B}}$ be a relation of $\mathbb{B}$ of some arity $r$, and let $\mathbb{G}$ be a classical gadget for $R^{\mathbb{B}}$ in $\mathbb{A}$, with distinguished vertices $g_{1},\dots,g_{r}$. Let $\mathbb{H}$ be a commutativity gadget for $\mathbb{A}$, with distinguished vertices $u,v$. We let $\mathbb{G}^{\prime}$ be the structure that initially consists of the disjoint union of $\mathbb{G}$ and a copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$ for each pair of distinct vertices $x,y$ in $G$, and in which we identify each such pair $x,y$ with the vertices $u,v$ from the copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$.
We show that $\mathbb{G}^{\prime}$ (with distinguished vertices $g_{1},\dots,g_{r}$) is a q-definition for $S$ in $\mathbb{A}$. Let $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ be a quantum homomorphism. In particular, $Q$ is non-contextual and can therefore be written as $h_{1}\oplus\dots\oplus h_{d}$ for some homomorphisms $h_{k}\colon\mathbb{R}\to\mathbb{A}_{S}$. Since $\mathbb{G}$ is a gadget for $S$ in $\mathbb{A}$, it follows that there exist homomorphisms $f_{k}\colon\mathbb{G}\to\mathbb{A}$ such that $f_{k}(g_{i})=h_{k}(i)$ for all $i\in[r]$. Then $Q^{\prime}=\bigoplus_{k\in[d]}f_{k}$ is a non-contextual quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$. In particular, $Q^{\prime}_{x}$ and $Q^{\prime}_{y}$ commute for all $x,y\in G$, and it follows from ($(\mathbf{c_{1}})$) that for all $x,y$ there exists a quantum homomorphism $Q^{(x,y)}\colon\mathbb{H}\to_{q}\mathbb{A}$ such that $Q^{(x,y)}_{u}=Q^{\prime}_{x}$ and $Q^{(x,y)}_{v}=Q^{\prime}_{y}$. We thus obtained a quantum homomorphism $\mathbb{G}^{\prime}\to_{q}\mathbb{A}$, which proves ($(\mathbf{q_{1}})$).
Let now $Q\colon\mathbb{G}^{\prime}\to_{q}\mathbb{A}$ be a quantum homomorphism. For each $x,y\in G$, we have by ($(\mathbf{c_{2}})$) that $Q_{x}$ and $Q_{y}$ commute. Thus, $Q$ restricts to a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ that is non-contextual and can thus be written as $h_{1}\oplus\dots\oplus h_{d}$, each $h_{k}$ being a classical homomorphism $\mathbb{G}\to\mathbb{A}$. Since $\mathbb{G}$ is a classical gadget for $R$, we obtain that $(h_{k}(g_{1}),\dots,h_{k}(g_{r}))\in R^{\mathbb{B}}$ for all $k\in[d]$. Thus, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}=\bigoplus_{k}(h_{k})_{g_{i},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$. ∎
Next, we show that q-definitions are compatible with quantum homomorphisms.
Proposition 17.
@@
Let $\mathbb{A},\mathbb{B}$ be structures, $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively. Then every quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is a quantum homomorphism $(A;R_{A})\to_{q}(B;R_{B})$.
@@
We prove $($$\mathbf{QH_{1}}$)). Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$. By ($(\mathbf{q_{1}})$), there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in\{1,\dots,r\}$. Let $S=Q\bullet Q^{\prime}$, which is a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{B}$. By ($(\mathbf{q_{2}})$), we have that the quantum function defined by $Q^{\prime\prime}_{i,b_{i}}:=S_{g_{i},b_{i}}$ for all $i\in\{1,\dots,r\}$ is a quantum homomorphism $Q^{\prime\prime}\colon\mathbb{R}\to(B,R_{B})$. Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$. Note that $S_{g_{i},b_{i}}$ is by definition $\sum_{a\in A}Q_{a,b_{i}}\otimes Q^{\prime}_{g_{i},a}=Q_{a_{i},b_{i}}\otimes\operatorname{id}$. Thus, $(\prod_{i=1}^{r}Q_{a_{i},b_{i}})\otimes\operatorname{id}=0$, from which we obtain that $\prod_{i=1}^{r}Q_{a_{i},b_{i}}=0$.

### v2

@@ 5 Quantum relational constructions
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $(A;S)$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
@@
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;S)$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
@@
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;S)$.
Thus, a q-definition is similar to a classical gadget where we impose (in ($(\mathbf{q_{2}})$)) the additional property that quantum homomorphisms $Q\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ must be non-contextual over the special vertices $\{g_{1},\dots,g_{r}\}$ of the gadget. In particular, in the case that $r=1$, then q-definitions and pp-definitions coincide.
Just as in the classical setting, we have a correspondence between the set of relations with a q-definition over a structure $\mathbb{A}$ and the set of relations $R$ such that $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(A;R)$.
Theorem 40.
Let $\mathbb{A}$ be a relational structure and let $R\subseteq A^{r}$. Then $R$ has a q-definition over $\mathbb{A}$ if, and only if, $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(A;R)$.
Proof.
Suppose first that $R$ has a q-definition over $\mathbb{A}$. Let $(\mathbb{G},g_{1},\dots,g_{r})$ be a q-definition, and pick a quantum polymorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ of some arity $n$. We need to show that $Q$ is an $n$-ary quantum polymorphism of $(A;R)$.
We first show $($$\mathbf{QH_{1}}$)). Let $\mathbf{a}^{1},\dots,\mathbf{a}^{n}\in R$ and $\mathbf{b}\not\in R$. For every $i\in[n]$, the classical map $h_{i}\colon[r]\to A$ defined by $j\mapsto a^{i}_{j}$ is a homomorphism $\mathbb{R}\to(A;R)$. By ($(\mathbf{q_{1}})$), there exists for all $i\in[n]$ a homomorphism $H_{i}\colon\mathbb{G}\to\mathbb{A}$ such that $H_{i}(g_{j})=h_{i}(j)$ for all $j\in[r]$. Define a quantum function $U\colon[r]\to_{\operatorname{q}}A$ by $U_{j,b}:=(Q\bullet(H_{1},\dots,H_{n}))_{g_{j},b}$. By ($(\mathbf{q_{2}})$), this is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;R)$ and therefore $\prod_{j=1}^{r}U_{j,b_{j}}=0$. But since $U_{j,b_{j}}=Q_{(a^{1}_{j},\dots,a^{n}_{j}),b_{j}}\otimes\operatorname{id}^{\otimes n}$, we obtain that $\prod_{j=1}^{r}Q_{(a^{1}_{j},\dots,a^{n}_{j}),b_{j}}=0$ and therefore $Q$ satisfies $($$\mathbf{QH_{1}}$)).
We now check $($$\mathbf{QH_{2}}$)). Let $\mathbf{a}^{1},\dots,\mathbf{a}^{n}\in R$, and let $i,j\in[r]$. Let $a,b\in A$. Using the same notation as above, we have $Q_{(a^{1}_{i},\dots,a^{n}_{i}),a},Q_{(a^{1}_{j},\dots,a^{n}_{j}),b}]\otimes\operatorname{id}^{\otimes n}=[U_{i,a},U_{j,b}]=0$.
We now prove the converse implication. Suppose that $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(A;R)$. Let $\mathbf{a}^{1},\dots,\mathbf{a}^{n}$ be an enumeration of all the tuples in $R$. We define the q-definition $\mathbb{G}$ to be the structure $\mathbb{A}^{n}$, and its $r$ distinguished vertices to be $(a^{1}_{i},\dots,a^{n}_{i})$ for $i\in[r]$.
We prove that ($(\mathbf{q_{2}})$) holds. Let $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ and define $Q^{\prime}_{i,b}=Q_{(r^{1}_{i},\dots,r^{n}_{i}),b}$, which is a quantum function $r]\to A$. Let $(b_{1},\dots,b_{r})\not\in R$. Then $\prod_{i=1}^{r}Q^{\prime}_{i,b_{i}}=\prod_{i=1}^{r}Q_{(r^{1}_{i},\dots,r^{n}_{i}),b_{i}}=0$ since $Q\in\operatorname{qPol}(A;R)$ by assumption and by $($$\mathbf{QH_{1}}$)). Thus, $($$\mathbf{QH_{1}}$)) holds for $Q^{\prime}$.
Similarly, if $i,j\in[r]$, $a,b\in A$, then $Q^{\prime}_{i,a},Q^{\prime}_{j,b}]=[Q_{(a^{1}_{i},\dots,a^{n}_{i}),a},Q_{(a^{1}_{j},\dots,a^{n}_{j}),b}]=0$, since $($$\mathbf{QH_{2}}$)) holds for $Q$ as a quantum polymorphism of $(A;R)$.
Finally, we check that ($(\mathbf{q_{1}})$) holds. Let $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;R)$ be a quantum homomorphism. Since $\mathbb{R}$ consists of a single tuple and $Q$ satisfies $($$\mathbf{QH_{2}}$)), one can write $Q$ as $\oplus_{i=1}^{d}h_{i}$, where each $h_{i}$ is a homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;R)$. For each $i$, there exists $i(j)\in[n]$ such that $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{i(j)}$. Define $g_{i}\colon\mathbb{A}^{n}\to\mathbb{A}$ to be the $n$-ary projection $\pi_{i(j)}$, and define $Q^{\prime}\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ to be $\oplus g_{i}$, concluding the proof of ($(\mathbf{q_{1}})$). ∎
@@
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{A}$ q-defines $\mathbb{B}$ if there exists a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
As an immediate corollary of Theorem 40, we obtain the following.
See 5
Proof.
It is enough to apply Theorem 40 to each of the relations of $\mathbb{B}$. ∎
Classical gadgets for CSPs that can be defined one from the other via pp-definitions provide logspace reductions between CSPs of different languages. It follows from the results above that the same holds for q-definitions.
@@
Let $\mathbb{A}$ and $\mathbb{B}$ be structures such that $\mathbb{A}$ q-defines $\mathbb{B}$. Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace.
Proof.
Applying Theorem 5, we know that $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(\mathbb{B})$. In particular, the inclusion map gives a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$. Then, the result follows from Theorem 3. ∎
Next, we show that q-definitions are compatible with quantum homomorphisms, in the following sense.
@@
Let $\mathbb{A},\mathbb{B}$ be structures, and let $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively. Then every quantum homomorphism $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ is a quantum homomorphism $(A;R_{A})\to_{\operatorname{q}}(B;R_{B})$.
@@
We prove $($$\mathbf{QH_{1}}$)). Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$. By ($(\mathbf{q_{1}})$), there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in[r]$. Let $S=Q\bullet Q^{\prime}$, which is a quantum homomorphism $\mathbb{G}\to_{\operatorname{q}}\mathbb{B}$. By ($(\mathbf{q_{2}})$), we have that the quantum function $Q^{\prime\prime}$ defined by $Q^{\prime\prime}_{i,b}:=S_{g_{i},b}$ for all $i\in[r]$ and all $b\in B$ is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(B;R_{B})$. Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$. Note that $S_{g_{i},b_{i}}$ is by definition $\sum_{a\in A}Q_{a,b_{i}}\otimes Q^{\prime}_{g_{i},a}=Q_{a_{i},b_{i}}\otimes\operatorname{id}$. Thus, $(\prod_{i=1}^{r}Q_{a_{i},b_{i}})\otimes\operatorname{id}=0$, from which we obtain that $\prod_{i=1}^{r}Q_{a_{i},b_{i}}=0$.